OpenStudy (anonymous):

will some one please help me with mixed fractions pleasee i really need help i will give medals!??

OpenStudy (ddcamp):

Sure, I can help.

OpenStudy (anonymous):

okay thanks so much its only five questions

OpenStudy (anonymous):

what is 7 times 2 3/7

OpenStudy (anonymous):

7 6/7 14 3/7 14 6/7 17

OpenStudy (ddcamp):

Remember that \[2\frac{ 3 }{ 7 } = 2+\frac{ 3 }{ 7 }\] Another way of writing 2 is 14/7 (because 14 divided by 7 = 2), so we have \[2+\frac{ 3 }{ 7 } = \frac{ 14 }{ 7 } + \frac{ 3 }{ 7 }= \frac{ 17 }{ 7 }\] So, what is 7 times 17/7?

OpenStudy (anonymous):

would i just multiply straight across???

OpenStudy (ddcamp):

You could. But remember, fractions are division So what this is saying is we're multiplying 17 by 7, then dividing by 7. What happens if you multiply then divide by the same number?

OpenStudy (anonymous):

no u do not

OpenStudy (anonymous):

so its 119/49???

OpenStudy (ddcamp):

No, 119/7; You only multiply the top part, not the bottom part

OpenStudy (anonymous):

okay if u give me the answers i will give u all my love and affection pleaseee

OpenStudy (anonymous):

???

OpenStudy (ddcamp):

This isn't an answers service. We;re here to help you learn to do this on your own.

OpenStudy (ddcamp):

\[7 \times \frac{ 17 }{ 7 } = \frac{ 119 }{ 7 } = 119 \div 7 = 17\]